# On the Robustness of Kernel Goodness-of-Fit Tests

### v3

@@ A.5.2 Proof of Theorem 4
$\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n})\;>\;q_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}\;$ $=\;\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\max(0,D(\mathbb{X}_{n})-\theta)\;>\;q_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}$
@@
where the first equality holds by Lemma 17 and noting that $D(\mathbb{X}_{n})$ is equivalent to the KSD between the empirical measure based on $\mathbb{X}_{n}$ and $P$, and the last line holds since $Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$. To show the first claim of our theorem, it suffices to show that the RHS of (51) converges to $\alpha$ as $n\to\infty$. Defining the term
@@
$\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\mathbb{S}_{P}(\mathbb{X}_{n},Q)\;>\;q_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}\;$ $=\;\Pr\nolimits_{\mathbf{W}}\big{(}D_{\mathbf{W}}(\mathbb{X}_{n})\;>\;q_{\infty,1-\alpha}(\mathbb{X}_{n})\;|\;\mathbb{X}_{\infty}\big{)}+\delta_{n}\;=\;\alpha+\delta_{n}\;,$
where the last step holds since $q_{\infty,1-\alpha}(\mathbb{X}_{n})$ is the $(1-\alpha)$-quantile of the conditional distribution of $D_{\mathbf{W}}(\mathbb{X}_{n})$ given $\mathbb{X}_{\infty}$. Moreover, since we have shown that $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})^{2}]<\infty$, we can apply Lemma 20 to conclude that $\delta_{n}\to 0$ as $n\to\infty$. Hence, the probability on the LHS of the above equation converges to $\alpha$. Combining with (51), we have proved that $\lim_{n\to\infty}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n})\;>\;q_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}\leq\alpha$. Taking supremum over all $Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)\cap\mathcal{P}(\mathbb{R}^{d};w)$ shows the first claim of our theorem.
@@
$\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim Q}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})^{2}]\;\leq\;\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})u_{p}(\mathbf{X}^{\prime},\mathbf{X}^{\prime})]\;$ $=\;\big{(}\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})]\big{)}^{2}\;\leq\;\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})^{2}]\;,$
@@
[…]
Combing these arguments shows that in the probability in (53), the LHS converges to a non-degenerate distribution, while the RHS tends to $\infty$- ∞. This implies (53) converges to 1, thus proving the second claim.
@@ A.6 Connections Between KSD Balls and Contamination Models
This section provides proofs for the results in Section 4.2.
A.6.1 Statement and Proof of Lemma 21
@@
Proof of Lemma 21  The assumption $\mathbb{E}_{\mathbf{Y}\sim R}[u_{p}(\mathbf{Y},\mathbf{Y})^{1/2}]<\infty$ implies that $D(R,P)$ is well-defined by Gorham and Mackey, (2017, Proposition 2). Moreover, when $k\in\mathcal{C}^{(1,1)}_{b}$ and $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$, Gorham and Mackey, (2015, Proposition 1) shows that $\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\cdot)]=0$. Using this and the linearity of expectation,
$D^{2}(Q,P)\;$ $=\;(1-\epsilon)^{2}\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim P}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})]+2(1-\epsilon)\epsilon\mathbb{E}_{\mathbf{X}\sim P,\mathbf{Y}\sim R}[u_{p}(\mathbf{X},\mathbf{Y})]+\epsilon^{2}\mathbb{E}_{\mathbf{Y},\mathbf{Y}^{\prime}\sim R}[u_{p}(\mathbf{Y},\mathbf{Y}^{\prime})]$
@@
A.6.2 Proof of Proposition 5
The proof follows a similar approach as Chérief-Abdellatif and Alquier, (2022, Lemma 3.3). Under the assumed kernel conditions, Lemma 2 shows that $\sup_{\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})\leq\tau_{\infty}<\infty$, which implies $D^{2}(Q,P)=\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim Q}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})]\leq\tau_{\infty}<\infty$ for any probability measure $Q\in\mathcal{P}(\mathbb{R}^{d})$. In particular, $D(Q,P)$ is well-defined for any $Q\in\mathcal{P}(\mathbb{R}^{d})$ and the conditions in Lemma 21 are met. For any $Q=(1-\epsilon)P+\epsilon R$ with $Q\in\mathcal{P}(P;\epsilon_{0})$, applying Lemma 21 shows that
@@ A.6.2 Proof of Proposition 5
$\sup_{Q\in\mathcal{P}(P;\epsilon_{0})}D(Q,P)\;\geq\;\sup_{\mathbf{z}\in\mathbb{R}^{d},\epsilon\in[0,\epsilon_{0}]}D\big{(}(1-\epsilon)P+\epsilon\delta_{\mathbf{z}},P\big{)}\;=\;\sup_{\mathbf{z}\in\mathbb{R}^{d},\epsilon\in[0,\epsilon_{0}]}\epsilon D(\delta_{\mathbf{z}},P)\;$ $=\;\epsilon_{0}\sup_{\mathbf{z}\in\mathbb{R}^{d}}u_{p}(\mathbf{z},\mathbf{z})^{\frac{1}{2}}$
@@
where the first inequality holds since $R=\delta_{\mathbf{z}}$ is only one possible type of perturbation, the first equality follows from Lemma 21, the second equality holds because the supremum over $\epsilon$ is reached at $\epsilon_{0}$, and the last step holds as $\sup_{\mathbf{z}\in\mathbb{R}^{d}}u_{p}(\mathbf{z},\mathbf{z})=\tau_{\infty}$ by definition. Combining with the upper bound (54), we have shown that $\sup_{Q\in\mathcal{P}(P;\epsilon_{0})}D(Q,P)=\epsilon_{0}\tau_{\infty}^{1/2}$.
A.6.3 Proof of Proposition 6
Under the assumed conditions on $k$, Lemma 2 shows that the Stein kernel is bounded by $\tau_{\infty}=\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})<\infty$, so in particular $D(Q,P)$ is well-defined for all $Q\in\mathcal{P}(\mathbb{R}^{d})$. Under the assumed integrability condition, we have $\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\cdot)]=0$ as argued in the paragraph before (11). We can therefore rewrite the squared KSD as
$D^{2}(Q,P)\;$ $=\;\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})q(\mathbf{x})q(\mathbf{x}^{\prime})\mathop{}\!\mathrm{d}\mathbf{x}\mathop{}\!\mathrm{d}\mathbf{x}^{\prime}\;=\;\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})(q(\mathbf{x})-p(\mathbf{x}))(q(\mathbf{x}^{\prime})-p(\mathbf{x}^{\prime}))\mathop{}\!\mathrm{d}\mathbf{x}\mathop{}\!\mathrm{d}\mathbf{x}^{\prime}\;,$
Using the assumed bound $|q(\mathbf{x})-p(\mathbf{x})|\leq\delta(\mathbf{x})$ and Lemma 2, we can bound the RHS of the above line by
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})\delta(\mathbf{x})\delta(\mathbf{x}^{\prime})\mathop{}\!\mathrm{d}\mathbf{x}\mathop{}\!\mathrm{d}\mathbf{x}^{\prime}\;\leq\;\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}|u_{p}(\mathbf{x},\mathbf{x}^{\prime})|\delta(\mathbf{x})\delta(\mathbf{x}^{\prime})\mathop{}\!\mathrm{d}\mathbf{x}\mathop{}\!\mathrm{d}\mathbf{x}^{\prime}\;$ $\leq\;\tau_{\infty}\bigg{(}\int_{\mathbb{R}^{d}}\delta(\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}\bigg{)}^{2}\;=\;\tau_{\infty}\delta_{0}^{2}\;.$
@@ A.6.4 KSD Balls and Fat Tails
The next proposition serves as an example for how Proposition 6 can be applied to design a robust KSD test when the model is Gaussian but data are drawn from t-distributions that are moment-matched to the model.
@@
Let $P=\mathcal{N}(0,1)$ and let $Q_{\nu}=t_{\nu}\sqrt{(\nu-2)/\nu}$, where $t_{\nu}$ is the t-distribution with degree-of-freedom (dof) $\nu>2$. Denote by $p$ and $q_{\nu}$ their probability density functions, and by $F_{\infty}$ and $F_{\nu}$ their cumulative distribution functions, respectively. Then $p(x)$ and $q_{\nu}(x)$ have exactly two intersections $a_{1}<a_{2}$ on $(0,\infty)$. If furthermore $k$ satisfies the conditions in Lemma 2, then $\mathrm{KSD}(Q_{\nu},P)\leq 4\tau_{\infty}^{1/2}(F_{\nu}(a_{1})-F_{\infty}(a_{1})+F_{\infty}(a_{2})-F_{\nu}(a_{2}))$.
This result suggests that, to ensure quantitative robustness against the scaled t-distribution with $\nu$ degrees-of-freedom, we can choose the uncertainty radius in the robust-KSD test to be $\theta=4\tau_{\infty}^{1/2}(F_{\nu}(a_{1})-F_{\infty}(a_{1})+F_{\infty}(a_{2})-F_{\nu}(a_{2}))$. The intersection points $a_{1},a_{2}$ do not have a closed form; instead, we approximate them by numerical solvers, which is trivial for this one-dimensional problem, and the numerical error is negligible as evidenced empirically in Figure 4.
Proof of Proposition 22  For $\nu>2$, the change-of-variable formula shows that the scaled t-distribution $Q_{\nu}$ has density function $q(x)=Z_{\nu}q_{\nu}^{\ast}(x)$, where $Z_{\nu}\coloneqq\Gamma(\frac{\nu+1}{2})/\big{(}\sqrt{\pi(\nu-2)}\Gamma(\frac{\nu}{2})\big{)}$, $q_{\nu}^{\ast}(x)\coloneqq(1+\frac{x^{2}}{\nu-2})^{-(\nu+1)/2}$, and $\Gamma(\cdot)$ is the Gamma function (see, e.g., Forbes et al.,, 2011, Chapter 8.1). To show $q_{\nu}(x)$ intersects with $p(x)$ at exactly two points on $(0,\infty)$, we first show that $q_{\nu}(x)$ is strong super-Gaussian (Palmer et al.,, 2010), i.e., $x\mapsto\log q_{\nu}(\sqrt{x})$ is convex on $[0,\infty)$. This is immediate by writing
@@
and noting that $u\mapsto\log(1+u)$ is concave on $[0,\infty)$. Palmer et al., (2010, Theorem 2) shows that any symmetric and strongly super-Gaussian density belongs to the class of density cross inequalities ($DC_{+}$), which are symmetric densities that cross a Gaussian density of equal variance exactly four times on $\mathbb{R}$, and take higher values than that normal density at $x=0$ and as $x\to\infty$. Since in this case both $Q_{\nu}$ and $P$ have unit variance, we conclude that $q_{\nu}$ and $p$ intersect at exactly four points on $\mathbb{R}$. By symmetry of these densities about $x=0$, this shows that $q_{\nu}$ and $p$ must intersect at exactly two points on $(0,\infty)$.
Call these two intersections $a_{1},a_{2}$ and assume without loss of generality $0<a_{1}<a_{2}$. It then follows that $q_{\nu}(x)\geq p(x)$ on $[0,a_{1})\cup[a_{2},\infty)$ and $q_{\nu}(x)\leq p(x)$ on $[a_{1},a_{2})$. We therefore have
@@
where in the second line we have again used the symmetry of normal and t densities. Applying Proposition 6 completes the proof.

### v4

@@ A.5.2 Proof of Theorem˜4
$\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n})\;>\;q_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}$
$=\;\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\max(0,D(\mathbb{X}_{n})-\theta)\;>\;q_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}$
@@
where the first equality holds by Lemma˜17 and noting that $D(\mathbb{X}_{n})$ is equivalent to the KSD between the empirical measure based on $\mathbb{X}_{n}$ and $P$, and the last line holds since $Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)$. To show the first claim of our theorem, it suffices to show that the RHS of (51) converges to $\alpha$ as $n\to\infty$. Defining the following bootstrapping error
@@
$\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\mathbb{S}_{P}(\mathbb{X}_{n},Q)\;>\;q_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}\;$ $=\;\Pr\nolimits_{\mathbf{W}}\big{(}D_{\mathbf{W}}(\mathbb{X}_{n})\;>\;q_{\infty,1-\alpha}(\mathbb{X}_{n})\;|\;\mathbb{X}_{\infty}\big{)}+\delta_{n}$
$=\;\alpha+\delta_{n}\;,$ (52)
where the last step holds since $q_{\infty,1-\alpha}(\mathbb{X}_{n})$ is the $(1-\alpha)$-quantile of the conditional distribution of $D_{\mathbf{W}}(\mathbb{X}_{n})$ given $\mathbb{X}_{\infty}$. Moreover, since we have shown that $\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})^{2}]<\infty$, we can apply Lemma˜20 to conclude that $\delta_{n}\to 0$ as $n\to\infty$. Hence, the probability on the LHS of the above equation converges to $\alpha$. Combining with (51), we have proved that $\lim_{n\to\infty}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n})\;>\;q_{\infty,1-\alpha}(\mathbb{X}_{n})\big{)}\leq\alpha$. Taking supremum over all $Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)\cap\mathcal{P}(\mathbb{R}^{d};w)$ shows the first claim of our theorem.
@@
$\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim Q}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})^{2}]\;\leq\;\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})u_{p}(\mathbf{X}^{\prime},\mathbf{X}^{\prime})]\;$ $=\;\big{(}\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})]\big{)}^{2}$
$\leq\;\mathbb{E}_{\mathbf{X}\sim Q}[u_{p}(\mathbf{X},\mathbf{X})^{2}]\;,$
@@
$\lim_{n\to\infty}(q_{\infty,1-\alpha}(\mathbb{X}_{n})+\theta)^{2}-D^{2}(Q,P)\;=\;\theta^{2}-D^{2}(Q,P)\;<\;0\;.$
Combing these arguments shows that in the probability in (54), the LHS converges to a non-degenerate distribution, while the RHS tends to -$\infty$. This implies (54) converges to 1, thus proving the second claim.
Remark 20.1.
If the bootstrapped quantile $q_{\infty,1-\alpha}(\mathbb{X}_{n})$ is replaced by the quantile $q^{\ast}_{1-\alpha}(\mathbb{X}_{n})$ of the distribution of the $P$-KSD $\mathbb{S}_{P}(\mathbb{X}_{n},Q)$, then the bootstrap approximation error in (52) vanishes, i.e., $\delta_{n}=0$ for all $n$. We can then prove the following stronger, uniform level control
$\limsup_{n\to\infty}\sup_{Q\in\mathcal{B}^{\mathrm{KSD}}(P;\theta)\cap\mathcal{P}(\mathbb{R}^{d};w)}\Pr\nolimits_{\mathbb{X}_{n}\sim Q,\mathbf{W}}\big{(}\Delta_{\theta}(\mathbb{X}_{n})\;>\;q^{\ast}_{1-\alpha}(\mathbb{X}_{n})\big{)}\;\leq\;\alpha\;.$
In fact, the inequality holds for any finite $n$, since the RHS of (52) equals to $\alpha$ for all $n$.
@@ A.6 Connections Between KSD Balls and Contamination Models
This section provides proofs for the results in Section˜4.2.
A.6.1 Statement and Proof of Lemma˜21
@@
Proof of Lemma˜21  The assumption $\mathbb{E}_{\mathbf{Y}\sim R}[u_{p}(\mathbf{Y},\mathbf{Y})^{1/2}]<\infty$ implies that $D(R,P)$ is well-defined by Gorham and Mackey, (2017, Proposition 2). Moreover, when $k\in\mathcal{C}^{(1,1)}_{b}$ and $\mathbb{E}_{\mathbf{X}\sim P}[\|\mathbf{s}_{p}(\mathbf{X})\|_{2}]<\infty$, Gorham and Mackey, (2015, Proposition 1) shows that $\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\cdot)]=0$. Using this and the linearity of expectation,
$D^{2}(Q,P)\;$ $=\;(1-\epsilon)^{2}\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim P}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})]+2(1-\epsilon)\epsilon\mathbb{E}_{\mathbf{X}\sim P,\mathbf{Y}\sim R}[u_{p}(\mathbf{X},\mathbf{Y})]$
$\;\quad+\epsilon^{2}\mathbb{E}_{\mathbf{Y},\mathbf{Y}^{\prime}\sim R}[u_{p}(\mathbf{Y},\mathbf{Y}^{\prime})]$
@@
A.6.2 Proof of Proposition˜5
The proof follows a similar approach as Chérief-Abdellatif and Alquier, (2022, Lemma 3.3). Under the assumed kernel conditions, Lemma˜2 shows that $\sup_{\mathbf{x},\mathbf{x}^{\prime}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})\leq\tau_{\infty}<\infty$, which implies $D^{2}(Q,P)=\mathbb{E}_{\mathbf{X},\mathbf{X}^{\prime}\sim Q}[u_{p}(\mathbf{X},\mathbf{X}^{\prime})]\leq\tau_{\infty}<\infty$ for any probability measure $Q\in\mathcal{P}(\mathbb{R}^{d})$. In particular, $D(Q,P)$ is well-defined for any $Q\in\mathcal{P}(\mathbb{R}^{d})$ and the conditions in Lemma˜21 are met. For any $Q=(1-\epsilon)P+\epsilon R$ with $Q\in\mathcal{P}(P;\epsilon_{0})$, applying Lemma˜21 shows that
@@ A.6.2 Proof of Proposition˜5
$\sup_{Q\in\mathcal{P}(P;\epsilon_{0})}D(Q,P)\;\geq\;\sup_{\mathbf{z}\in\mathbb{R}^{d},\epsilon\in[0,\epsilon_{0}]}D\big{(}(1-\epsilon)P+\epsilon\delta_{\mathbf{z}},P\big{)}\;$ $=\;\sup_{\mathbf{z}\in\mathbb{R}^{d},\epsilon\in[0,\epsilon_{0}]}\epsilon D(\delta_{\mathbf{z}},P)$
$=\;\epsilon_{0}\sup_{\mathbf{z}\in\mathbb{R}^{d}}u_{p}(\mathbf{z},\mathbf{z})^{\frac{1}{2}}$
@@
where the first inequality holds since $R=\delta_{\mathbf{z}}$ is only one possible type of perturbation, the first equality follows from Lemma˜21, the second equality holds because the supremum over $\epsilon$ is reached at $\epsilon_{0}$, and the last step holds as $\sup_{\mathbf{z}\in\mathbb{R}^{d}}u_{p}(\mathbf{z},\mathbf{z})=\tau_{\infty}$ by definition. Combining with the upper bound (55), we have shown that $\sup_{Q\in\mathcal{P}(P;\epsilon_{0})}D(Q,P)=\epsilon_{0}\tau_{\infty}^{1/2}$.
A.6.3 Proof of Proposition˜6
Under the assumed conditions on $k$, Lemma˜2 shows that the Stein kernel is bounded by $\tau_{\infty}=\sup_{\mathbf{x}\in\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x})<\infty$, so in particular $D(Q,P)$ is well-defined for all $Q\in\mathcal{P}(\mathbb{R}^{d})$. Under the assumed integrability condition, we have $\mathbb{E}_{\mathbf{X}\sim P}[u_{p}(\mathbf{X},\cdot)]=0$ as argued in the paragraph before (11). We can therefore rewrite the squared KSD as
$D^{2}(Q,P)\;$ $=\;\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})q(\mathbf{x})q(\mathbf{x}^{\prime})\mathop{}\!\mathrm{d}\mathbf{x}\mathop{}\!\mathrm{d}\mathbf{x}^{\prime}$
$=\;\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})(q(\mathbf{x})-p(\mathbf{x}))(q(\mathbf{x}^{\prime})-p(\mathbf{x}^{\prime}))\mathop{}\!\mathrm{d}\mathbf{x}\mathop{}\!\mathrm{d}\mathbf{x}^{\prime}\;,$
Using the assumed bound $|q(\mathbf{x})-p(\mathbf{x})|\leq\delta(\mathbf{x})$ and Lemma˜2, we can bound the RHS of the above line by
$\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}u_{p}(\mathbf{x},\mathbf{x}^{\prime})\delta(\mathbf{x})\delta(\mathbf{x}^{\prime})\mathop{}\!\mathrm{d}\mathbf{x}\mathop{}\!\mathrm{d}\mathbf{x}^{\prime}\;$ $\leq\;\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}|u_{p}(\mathbf{x},\mathbf{x}^{\prime})|\delta(\mathbf{x})\delta(\mathbf{x}^{\prime})\mathop{}\!\mathrm{d}\mathbf{x}\mathop{}\!\mathrm{d}\mathbf{x}^{\prime}$
$\leq\;\tau_{\infty}\bigg{(}\int_{\mathbb{R}^{d}}\delta(\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}\bigg{)}^{2}\;=\;\tau_{\infty}\delta_{0}^{2}\;.$
@@ A.6.4 KSD Balls and Fat Tails
The next proposition serves as an example for how Proposition˜6 can be applied to design a robust KSD test when the model is Gaussian but data are drawn from t-distributions that are moment-matched to the model.
@@
Let $P=\mathcal{N}(0,1)$ and let $Q_{\nu}=t_{\nu}\sqrt{(\nu-2)/\nu}$, where $t_{\nu}$ is the t-distribution with degree-of-freedom (dof) $\nu>2$. Denote by $p$ and $q_{\nu}$ their probability density functions, and by $F_{\infty}$ and $F_{\nu}$ their cumulative distribution functions, respectively. Then $p(x)$ and $q_{\nu}(x)$ have exactly two intersections $a_{1}<a_{2}$ on $(0,\infty)$. If furthermore $k$ satisfies the conditions in Lemma˜2, then $\mathrm{KSD}(Q_{\nu},P)\leq 4\tau_{\infty}^{1/2}(F_{\nu}(a_{1})-F_{\infty}(a_{1})+F_{\infty}(a_{2})-F_{\nu}(a_{2}))$.
This result suggests that, to ensure quantitative robustness against the scaled t-distribution with $\nu$ degrees-of-freedom, we can choose the uncertainty radius in the robust-KSD test to be $\theta=4\tau_{\infty}^{1/2}(F_{\nu}(a_{1})-F_{\infty}(a_{1})+F_{\infty}(a_{2})-F_{\nu}(a_{2}))$. The intersection points $a_{1},a_{2}$ do not have a closed form; instead, we approximate them by numerical solvers, which is trivial for this one-dimensional problem, and the numerical error is negligible as evidenced empirically in Figure˜4.
Proof of Proposition˜22  For $\nu>2$, the change-of-variable formula shows that the scaled t-distribution $Q_{\nu}$ has density function $q(x)=Z_{\nu}q_{\nu}^{\ast}(x)$, where $Z_{\nu}\coloneqq\Gamma(\frac{\nu+1}{2})/\big{(}\sqrt{\pi(\nu-2)}\Gamma(\frac{\nu}{2})\big{)}$, $q_{\nu}^{\ast}(x)\coloneqq(1+\frac{x^{2}}{\nu-2})^{-(\nu+1)/2}$, and $\Gamma(\cdot)$ is the Gamma function (see, e.g., Forbes et al.,, 2011, Chapter 8.1). To show $q_{\nu}(x)$ intersects with $p(x)$ at exactly two points on $(0,\infty)$, we first show that $q_{\nu}(x)$ is strong super-Gaussian (Palmer et al.,, 2010), i.e., $x\mapsto\log q_{\nu}(\sqrt{x})$ is convex on $0,\infty)$. This is immediate by writing
@@
and noting that $u\mapsto\log(1+u)$ is concave on $0,\infty)$. Palmer et al., (2010, Theorem 2) shows that any symmetric and strongly super-Gaussian density belongs to the class of density cross inequalities ($DC_{+}$), which are symmetric densities that cross a Gaussian density of equal variance exactly four times on $\mathbb{R}$, and take higher values than that normal density at $x=0$ and as $x\to\infty$. Since in this case both $Q_{\nu}$ and $P$ have unit variance, we conclude that $q_{\nu}$ and $p$ intersect at exactly four points on $\mathbb{R}$. By symmetry of these densities about $x=0$, this shows that $q_{\nu}$ and $p$ must intersect at exactly two points on $(0,\infty)$.
Call these two intersections $a_{1},a_{2}$ and assume without loss of generality $0<a_{1}<a_{2}$. It then follows that $q_{\nu}(x)\geq p(x)$ on $0,a_{1})\cup[a_{2},\infty)$ and $q_{\nu}(x)\leq p(x)$ on $a_{1},a_{2})$. We therefore have
@@
where in the second line we have again used the symmetry of normal and t densities. Applying Proposition˜6 completes the proof.